# GENESIS-RL: GEnerating Natural Edge-cases with Systematic Integration of Safety considerations and Reinforcement Learning

### v1

@@ V-B2 Initial scene generator
Scenic [30] was utilized for creating the initial scenes as a strategic approach to ensure precision and versatility in our experimental setup. It allows us to craft realistic, detailed and specific initial scenes, thereby enhancing the relevance and challenge of each test instance presented to the system. Specifically, this generator is tasked with assigning the positions of vehicles within the simulator: For each initial scene, we introduce 30 vehicles, including the ego vehicle, a lead vehicle, a neighboring vehicle on the front right, and the remaining 27 vehicles randomly positioned within a specific radius of the ego vehicle. Vehicle colors, models, and the distance between the ego, lead, and neighbor vehicles could be either varied randomly or deterministic, depending on the use case.
We leveraged CARLA’s [18] record feature to efficiently catalog the configurations of the initial scenes. These saved scenes are then randomly sampled and loaded back into the simulation between training or testing episodes. A total of 1199 initial scenes for Town05 and 1202 for Town10 were generated and cataloged using the this generator.
@@ V-B4 Reward calculator
The reward $r_{t}$ is calculated by the weighted sum of the learning module loss $r_{m}$ and the violation score $r_{v}$. The learning module loss is defined by the Intersection Over Union (IOU) metric, assessing the system’s object detection precision in real-time. Simultaneously, the violation score is generated by the rulebook evaluating a trajectory, focusing on two critical safety rules:
  • Vehicle collision, violated if the ego vehicle collides with other vehicles.
  • Vehicle proximity, violated when the ego vehicle approaches closer than a predetermined minimum distance to the vehicle in front.
The reward $r_{t}$ is and calculated as follows:
$r_{t}$ $=r_{m}+r_{v}$ (1)
$r_{i}$ $=e^{-iou}$ (2)
$r_{v}$ $=\sum_{k}ln(1+w_{k}s_{k}),\ k=c,p$ (3)
where $iou$ is the IOU metric, $c,p$ represent collision and proximity rules, respectively, $(w_{c},w_{p})=(500,100)$, and $s_{k}$ is the speed of the ego vehicle in meters per second when a violation occurs and serves as a measure to quantify the resultant impact incurred by the violation. The predetermined minimum distance for the proximity rule is set to 5 meters.
@@ V-C Training and Testing
Our experiments are divided into two phases: training and testing, each aimed at evaluating the effectiveness of the GENESIS-RL framework.
@@ V-C1 Training
The training experiments was conducted in CARLA’s  [18] Town10, an urban map setting with multiple four-way intersections. To ensure the DRL agent experiences a broad range of initial scene configurations, we start each episode by randomly selecting from 1202 pre-cataloged scenes, each featuring vehicles with randomized colors and makes. This setup is further enhanced by the stochastic behaviors of non-ego vehicles, governed by CARLA’s [18] traffic manager, introducing both realism and complexity. The variable vehicle arrangement, appearances, and the other vehicle agents’ stochastic nature significantly enriched the training landscape, equipping the DRL agent to adeptly uncover system’s vulnerabilities. The DRL agent is trained with 40960 steps, with an episode length of 512. The policy of the DRL agent updates every four trajectories.
@@ V-C2 Testing
Detailed records of violation scores and other critical data were kept, with averages calculated for thorough analysis. These findings were benchmarked against two specific scenarios: our system navigating in clear weather and under randomly perturbed weather conditions by a non-strategic agent, providing a comprehensive evaluation of the GENESIS-RL framework. The length of each testing episode is identical that of the training, which is 512 time steps in the simulation.
@@ VI Results and Discussion
In this section, we present the outcomes of our experiments, systematically analyzing the performance and efficacy of the GENESIS-RL framework. The results underscore the capability of our approach to generate meaningful edge cases, highlighting key findings and insights gained through testing scenarios. Particularly, we validate our approach by answering the following questions:
@@
To begin our exploration of these questions, we first examine the learning dynamics of GENESIS-RL, as evidenced by the progression of rewards over time. Fig. 2 presents the reward plot from five training runs, showcasing the DRL agent’s ability to adapt and improve through its interactions through time. Each run is depicted as a dashed line within the figure. The mean reward across these five runs is displayed as a solid blue line, indicating the overall learning trend. Additionally, the variance in rewards across the runs is represented by the shaded area, offering insight into the consistency and reliability of the learning process.
Figure 2: Reward plot for GENESIS-RL training runs. This figure illustrates the variation in rewards over training steps in five training runs. The solid blue line indicates the average reward of the runs, while the light blue shaded area represents the variance. Dashed lines show individual run trajectories, highlighting the spread around the average.
@@ VI-A1 Proximity violation score increase
Under weather scenarios controlled by GENESIS-RL, the system’s braking response was markedly delayed compared to its reaction under other testing scenarios and, in some instances, the braking behavior is altogether absent. This delay/absence is quantifiably demonstrated through the analysis of the ego vehicle’s telemetry data (See Fig. 4(c), the system operating under the GENESIS-RL policy maintains a brakes much later, i.e., the blue curve is shifted more to the right compared to the two other scenarios), showcasing a contrast in the system’s ability to maintain safe following distances under varied environmental influences.
@@ VI-A2 Collision violation score increase
  • Delayed detection collision: Occurrences where the system detects the lead vehicle too late to stop in time, leading to collisions at reduced speeds (See Fig. 4(c). These incidents highlights the deficiencies in the system’s detection mechanisms.
@@ VI-B Effectiveness of GENESIS-RL in Generating Edge Cases
Our results show that with the GENESIS-RL framework, we were able to generate edge cases that pose significant challenges not only to automated systems but also to human perception and response capabilities. A prime example of such conditions includes scenarios combining foggy weather with heavy rainfall, or nocturnal settings accentuated by heavy rain, where the reflections on wet surfaces severely disrupt the detection capabilities of autonomous systems. These conditions, inherently challenging due to their impact on visibility and sensor efficiency, highlight the scenario generation capabilities of GENESIS-RL, underscoring its potential for creating diverse testing environments that closely mimic real-world driving complexities.
During inference, we observed that the DRL agent relies heavily on manipulating certain parametric knobs (such as rain and fog density) to introduce challenges to the system. To investigate the capability of GENESIS-RL in generating versatile cases, we conducted experiments with fog and rain density levels set to zero, individually and in combination. Despite the absence of these parametric knobs, our framework is still capable of creating edge cases that undermine the system’s performance, demonstrating its robustness in edge case generation beyond the reliance on certain powerful and effective parametric knobs (See banner figure on first page for reference).
Figure 3: Violation scores and sum of minimum following distance deficit (based on RSS) across three testing scenarios - the system operates under sunny weather, random weather and under the GENESIS-RL policy.
Figure 4: Examples of the two system failure modes based on vehicle telemetry. (a) Example of a non-detection collision - where the system crashed into the vehicle in front at full/high speed due to the non-detection of the other vehicle. (b) Intermittent detection collision - where the intermittent detection prevents the ego vehicle from stopping in time, leading to a lower-speed collision. (c) Delayed detection collision - where the system detects the lead car too late to stop in time. In the tests depicted in this figure, the ego vehicle successfully avoided collisions across both sunny and random weather scenarios and only failed under the conditions generated by GENESIS-RL.
@@ VII Conclusion and Future Work
In conclusion, our study demonstrates the GENESIS-RL framework has demonstrated its capability to generate complex and challenging edge cases for autonomous systems, which are critical for thoroughly testing and enhancing the reliability of systems as such. Moreover, GENESIS-RL proved capable of producing edge cases that will lead to system failure even in the absence of some dominating factors, underscoring its robustness and potential broader application in safety-critical testing environments.
The implications of the results from our work is manifold. First, they highlight the need for including a wide range of challenging scenarios in the testing protocols for autonomous systems, ensuring they are well-equipped to handle intricacies of complex and dynamic environments. Additionally, the ability of GENESIS-RL to generate test conditions without relying exclusively on dominant factors showcases its value in crafting safer, more dependable autonomous system solutions.
Looking ahead, we plan to expand our exploration to a broader array of parameters, including a more extensive range of weather conditions, the behavior of other actors in the environment, and additional variables, to further enhance the scenario generation capabilities of GENESIS-RL and its application in creating even more diverse and challenging testing environments for autonomous systems.

### v2

@@ V-B2 Initial scene generator
Scenic [32] was utilized for creating the initial scenes as a strategic approach to ensure precision and versatility in our experimental setup. It allows us to craft realistic, detailed and specific initial scenes, thereby enhancing the relevance and challenge of each test instance presented to the system. Specifically, this generator is tasked with assigning the positions of vehicles within the simulator: For each initial scene, we introduce 30 vehicles, including the ego vehicle, a lead vehicle, a neighboring vehicle on the front right, and the remaining 27 vehicles randomly positioned within a specific radius of the ego vehicle. Vehicle colors, models, and the distance between the ego, lead, and neighbor vehicles could be either varied randomly or deterministic, depending on the use case. All vehicles except for the ego vehicle are set to CARLA’s autopilot in order to provide a dynamic environment.
We leveraged CARLA’s [18] record feature to efficiently catalog the configurations of the initial scenes into json files. These saved scenes are then randomly sampled and loaded back into the simulation between training or testing episodes. A total of 1199 initial scenes for Town05 and 1202 for Town10 were generated and cataloged using this generator.
@@ V-B4 Reward calculator
The reward $r_{t}$ is calculated by the weighted sum of the learning module loss $r_{m}$ and the violation score $r_{v}$. The learning module loss is defined by the Intersection Over Union (IOU) metric, assessing the system’s object detection precision in real-time. Simultaneously, the violation score is generated by the rulebook evaluating a trajectory $\tau=s_{1},s_{2},s_{3},....s_{T}$, consisting of the following two violation scores based on two critical safety rules:
  • Vehicle collision rule, $\lambda_{c}(\tau)$, violated if the ego vehicle collides with other vehicles:
    $\lambda_{c}(\tau)=\sum_{t=1}^{T}\alpha_{c}(s_{t})$ (1)
[…]
  • Vehicle proximity rule, $\lambda_{p}(\tau)$, violated when the distance $d$ between the ego vehicle and the vehicle in front is less than a predetermined distance (which is set to 5 meters):
    $\lambda_{p}(\tau)=\sum_{t=1}^{T}\alpha_{p}(s_{t})$ (3)
    $\centering\alpha_{p}(s_{t})=v_{p}\text{ if $d$ $<$ 5},0\text{ otherwise}\@add@centering$ (4)
where $v_{c}$ and $v_{p}$ are the speeds of the ego vehicle in meters per second when a violation occurs. $\alpha_{c}(s_{t})$ and $\alpha(s_{t})$ are then weighted by their respective weights $w_{c}$ and $w_{p}$ then summed and transformed using the natural logarithm to compute $r_{v}$. Finally, the reward $r_{t}$ is calculated as follows:
$r_{t}=r_{m}+r_{v}=e^{-iou}+ln(1+w_{c}\alpha_{c}+w_{p}\alpha_{p})$ (5)
where $iou$ is the IOU metric, $(w_{c},w_{p})=(500,100)$. The rationale for incorporating speed values when a violation occurs is to ensure that violations occurring at higher speeds are assigned higher violation scores.
@@ V-C Training and Testing
Our experiments are divided into two phases: training and testing, each aimed at evaluating the effectiveness of our framework.
@@ V-C1 Training
The training experiments were conducted in CARLA’s  [18] Town10, an urban map setting with multiple four-way intersections. To ensure the DRL agent experiences a broad range of initial scene configurations, we start each episode by randomly selecting from 1202 pre-cataloged scenes, each featuring vehicles with randomized colors and makes. This setup is further enhanced by the stochastic behaviors of non-ego vehicles, governed by CARLA’s [18] traffic manager, introducing both realism and complexity. The variable vehicle arrangement, appearances, and the other vehicle agents’ stochastic nature significantly enriched the training landscape, equipping the DRL agent to adeptly uncover the system’s vulnerabilities. The DRL agent is trained with 40960 steps, with an episode length of 512. The policy of the DRL agent updates every four trajectories.
@@ V-C2 Testing
Records of violation scores and other information were kept, with averages calculated for analysis. These findings were benchmarked against two specific scenarios: our system navigating in clear weather and under randomly perturbed weather conditions by a non-strategic agent, providing a comprehensive evaluation of the GENESIS-RL framework. The length of each testing episode is identical to that of the training, which is 512 time steps in the simulation.
@@ V-D Evaluation Metrics
V-E Computational Time
The experiments were conducted on a computational setup with a 32-core CPU, 64GB of RAM, and an NVIDIA GTX TITAN X graphics card. A complete training run required approximately 3-4 hours on this hardware, while a testing run took about 1-2 minutes per episode.
@@ VI Results and Discussion
In this section, we present the outcomes of our experiments, including an analysis of the performance and efficacy of our framework. The results underscore the capability of our approach to generate meaningful edge cases, highlighting key findings and insights gained through testing scenarios. We validate our approach by answering the following questions:
@@ VI-A1 Proximity violation score increase
Under weather scenarios controlled by GENESIS-RL, the system’s braking response was markedly delayed compared to its reaction under other testing scenarios and, in some instances, the braking behavior is altogether absent. This delay/absence is quantifiably demonstrated through the analysis of the ego vehicle’s telemetry data (See Fig. 3(c), the system operating under the GENESIS-RL policy maintains brakes much later, i.e., the blue curve is shifted more to the right compared to the two other scenarios), showcasing a contrast in the system’s ability to maintain safe following distances under varied environmental influences.
@@ VI-A2 Collision violation score increase
  • Delayed detection collision: Occurrences where the system detects the lead vehicle too late to stop in time, leading to collisions at reduced speeds (See Fig. 3(c). These incidents highlight the deficiencies in the system’s detection mechanisms.
@@ VI-B Effectiveness of GENESIS-RL in Generating Edge Cases
Our results show that with the GENESIS-RL framework, we were able to generate edge cases that pose significant challenges not only to automated systems but also to human perception and response capabilities. A prime example of such conditions includes scenarios combining foggy weather with heavy rainfall or nocturnal settings accentuated by heavy rain, where the reflections on wet surfaces severely disrupt the detection capabilities of autonomous systems. These conditions, inherently challenging due to their impact on visibility and sensor efficiency, highlight the scenario generation capabilities of GENESIS-RL, underscoring its potential for creating diverse testing environments that closely mimic real-world driving complexities.
During inference, we observed that the DRL agent relies heavily on manipulating certain parametric knobs (such as rain and fog density) to introduce challenges to the system. To investigate GENESIS-RL’s capability to generate a variety of edge cases, we conducted experiments with fog and rain density levels set to zero, individually and in combination. Despite the absence of these parametric knobs, our framework is still capable of creating edge cases that undermine the system’s performance, demonstrating its robustness in edge case generation beyond the reliance on certain powerful and effective parametric knobs (See banner figure on the first page for reference).
Figure 2: Violation scores and the sum of minimum following distance deficit (based on RSS) across three testing scenarios - the system operates under sunny weather, random weather and under the GENESIS-RL policy. The results presented are averages across 50 runs with randomly selected initial scenes.
Figure 3: Examples of the system failure modes based on vehicle telemetry. (a) Example of a non-detection collision - where the system crashed into the vehicle in front at full/high speed due to the non-detection of the other vehicle. (b) Intermittent detection collision - where the intermittent detection prevents the ego vehicle from stopping in time, leading to a lower-speed collision. (c) Delayed detection collision - where the system detects the lead car too late to stop in time. In the tests depicted in this figure, the ego vehicle successfully avoided collisions across both sunny and random weather scenarios and only failed under the conditions generated by GENESIS-RL.
@@ VII Conclusion and Future Work
In conclusion, our study demonstrates the GENESIS-RL framework’s capability to generate complex and challenging edge cases for autonomous systems, which are critical for thoroughly testing and enhancing the reliability of systems as such. Moreover, GENESIS-RL proved capable of producing edge cases that will lead to system failure even in the absence of some dominating factors, underscoring its robustness and potential broader application in safety-critical testing environments.
The implications of the results from our work is manifold. First, they highlight the need for including a wide range of challenging scenarios in the testing protocols for autonomous systems, ensuring they are well-equipped to handle the intricacies of complex and dynamic environments. Additionally, the ability of GENESIS-RL to generate test conditions without relying exclusively on dominant factors showcases its value in crafting safer, more dependable autonomous system solutions.
In future work, we aim to expand our exploration to include a wider spectrum of parameters, such as a more comprehensive range of weather parameters, behaviors of other actors (e.g., vehicles and pedestrians) to further enhance the scenario generation capabilities of GENESIS-RL, thereby facilitating the creation of more diverse and challenging testing environments for autonomous systems.